OpenStudy (anonymous):

Solve for x. |x – 2| + 10 = 12 x = 4 and x = 0 x = -15 and x = 9 x = -4 and x = -2 x = -3 and x = 1 no solutions @jiteshmeghwal9 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

Any ideas on the first step?

OpenStudy (anonymous):

x.Mod(x-2)=2

OpenStudy (jiteshmeghwal9):

x-2+10=12 =>x+8=12 =>x=12-8=4.

OpenStudy (anonymous):

no sulotion for this

OpenStudy (jiteshmeghwal9):

Don't ask same type of ques. again & again @kenzi.sloboda.14 .